OpenStudy (anonymous):

find the equation of a line that is perpendicular to the line 1/2x+2 and contains the points (5,0)

OpenStudy (anonymous):

slope of a line perpendicular to a line with slope m is -1/m so slope is -2 y=-2x+c c=10 y=-2x+10

OpenStudy (anonymous):

Problem: find the equation of a line that is perpendicular to the line 1/2x+2 and contains the points (5,0) Here's how I would approach it. When you have two perpendicular lines, if you take their two slopes and multiply them, you get -1. So the line that is given has a slope of 1/2. That means that the second line will have a slope of -2. So, I can write the equation so far as: y = -2x + b And b would be the y-intercept. I would then substitute in x and y values for the point known to lie on the second line. So I would have: 0 = -2(5) + b I'd solve for b and find that b = 10 So the full equation is: y = -2x + 10 Hope that helps!